# RefinedFields: Radiance Fields Refinement for Unconstrained Scenes

### v3

@@ Abstract
Modeling large scenes from unconstrained images has proven to be a major challenge in computer vision. Existing methods tackling in-the-wild scene modeling operate in closed-world settings, where no conditioning on priors acquired from real-world images is present. We propose RefinedFields, which is, to the best of our knowledge, the first method leveraging pre-trained models to improve in-the-wild scene modeling. We employ pre-trained networks to refine K-Planes representations via optimization guidance using an alternating training procedure. We carry out extensive experiments and verify the merit of our method on synthetic data and real tourism photo collections. RefinedFields enhances rendered scenes with richer details and improves upon its base representation on the task of novel view synthesis in the wild. Our project page can be found at https://refinedfields.github.io .
Computer Vision, 3D, Neural Radiance Fields, K-Planes, Tri-Planes, in the wild
@@ 1 Introduction
To draw novel objects and views, humans often rely on a blend of cognition and intuition, where the latter is built on a large prior acquired from a long-term continuous exploration of the visual world. Nevertheless, ablating one of these two elements results in catastrophic representations. On the one hand, humans find it particularly difficult to draw a bicycle based solely on this preconceived prior (Gimini, 2016). However, once one photograph is observed, drawing novel views of a bicycle becomes straightforward. On the other hand, drawing monuments and complex objects based solely on observed images, and with no preconceived notions of geometry and physics, is also non-trivial. In computer vision, recent methods tackling object generation and novel view synthesis are exclusively built on either the former or the latter ablation.
Refer to caption Refer to caption Refer to caption
@@
The first class of methods approaches novel view synthesis by learning scenes through rigorous cognition, as in dense observations of captured images. Although classic methodologies like structure-from-motion (Hartley & Zisserman, 2004) and image-based rendering (Shum et al., 2008) have previously tackled this problem, the field has recently seen substantial advancements thanks to neural representations coupled with volume rendering techniques (Mildenhall et al., 2020; Chan et al., 2022). Recent methods extend NeRFs (Martin-Brualla et al., 2021; Chen et al., 2022b; Yang et al., 2023) and Tri-Planes (Fridovich-Keil et al., 2023) to support learning from unconstrained “in-the-wild” photo collections by enabling robustness against illumination variations and transient occluders. These representations, however, do not learn any prior across scenes as they are trained from scratch for each scene. This means that these representations are learned in a closed-world setting, where the information scope is limited to the training set at hand.
@@
Table 1: Related work overview. RefinedFields leverages unconstrained tourist photo collections (Jin et al., 2020) and a pre-trained prior (Rombach et al., 2022) to refine K-Planes, our underlying scene representation. Due to size constraints, the referenced version of this table can be found in Appendix A.
@@
NeRF NeRF NVS
K-Planes K-Planes
NFD Tri-Planes OG
3DGen Tri-Planes
Latent-NeRF NeRF
DreamFusion NeRF
PixelNeRF NeRF NVS
NeRDi NeRF
RealFusion NeRF
DiffusioNeRF NeRF
NerfDiff Tri-Planes
3DiM
Zero-1-to-3
NeRF-W NeRF NVS-W
Ha-NeRF NeRF
CR-NeRF NeRF
RefinedFields K-Planes
The second class of methods tackles novel view synthesis and object generation by learning and leveraging priors over images and scenes, reminiscent of drawing from insights and intuition. These methods have also recently witnessed accelerated advancements. Current works leverage pre-trained networks to achieve Object Generation (OG) (Poole et al., 2023; Metzer et al., 2023) and Novel View Synthesis (NVS) (Yu et al., 2021; Jain et al., 2021; Liu et al., 2023; Melas-Kyriazi et al., 2023). Particularly, Liu et al. (2023) achieve NVS from single images only by simply fine-tuning a pre-trained latent diffusion model (Rombach et al., 2022). This proves pivotal significance related to large-scale pre-trained vision models, as it shows that, although trained on 2D data, these models learn a rich geometric 3D prior about the visual world. Nevertheless, as these pre-trained models alone have no explicit multi-view geometric constraints, their use for 3D applications is usually prone to geometric issues (e.g. geometric inconsistencies, multi-face Janus problem, content drift issues (Shi et al., 2023, Figure 1)). This class of methods has not yet been explored for in-the-wild scene modeling, as leveraging priors over representations modeling unconstrained scenes is not evident.
RefinedFields proposal. Our work builds on the previous discourse and aims to enhance in-the-wild scene modeling by leveraging pre-trained networks. We adopt K-Planes (Fridovich-Keil et al., 2023) as a base scene representation, notably for its planar structure that makes it compatible with image-based networks. RefinedFields refines scene representations by projecting them onto the space of representations inferable by a pre-trained network, which pushes K-Planes features to more closely resemble real-world images. To do so, we present an alternating training procedure that iteratively switches between optimizing a K-Planes representation on images from a particular dataset, and fine-tuning a pre-trained network to output a new conditioning leading to a refined version of this K-Planes representation. Overall, this procedure guides the optimization of a particular scene, by leveraging not only the training dataset at hand but also the rich prior lying within the weights of the pre-trained model, leading to a better representation of fine details in the rendered scene.
We carry out extensive experiments and conduct quantitative and qualitative evaluations of RefinedFields. We show that our method improves upon K-Planes by providing richer details in scene renderings. We prove via ablation studies of our proposed work that this added value indeed comes from the fine-tuned prior of the pre-trained network. Figure 1 illustrates the improvements our method showcases on the Trevi fountain scene from Phototourism (Jin et al., 2020).
A summary of our contributions can be found below.
  • We introduce RefinedFields, a novel way to refine scene representations. This is, to the best of our knowledge, the first method leveraging pre-trained networks for novel view synthesis in-the-wild.
  • We show that the proposed scene refining pipeline, coupled with our training procedure, makes RefinedFields improve upon its baseline on the task of novel view synthesis in-the-wild.
  • The code for RefinedFields will be publicly available as open-source.
2 Related Work
RefinedFields achieves geometrically consistent novel view synthesis in-the-wild by leveraging unconstrained image collections from Phototourism (Jin et al., 2020), and a large-scale pre-trained network. Our method is the first to satisfy all of these attributes, as summarized in Table 1 which presents an overview of recent methods. In this section, we develop the various preceding works from which our method takes inspiration.
Neural representations.
Neural rendering (Tewari et al., 2020) has seen significant advancements since the introduction of NeRF (Mildenhall et al., 2020). At its core, neural rendering blends approaches from classical computer graphics (Wood et al., 2002; Cohen & Szeliski, 2014; Waechter et al., 2014) and machine learning (Park et al., 2019; Genova et al., 2020; Niemeyer et al., 2020) to reconstruct scenes from real-world observations. NeRF learns a scene by overfitting the weights of a neural network on posed images of said scene. This subsequently enables the reconstruction of the scene thanks to volume rendering (Kajiya & Herzen, 1984). Chan et al. (2022) introduce the Tri-Planes representation as a middle ground between implicit and explicit representations, enabling faster learning of scenes. More recently, Kerbl et al. (2023) tackle real-time high-quality radiance field rendering on unbounded scenes. Note that rendering unbounded scenes is dissimilar to our task of in-the-wild rendering, where the latter consists of learning scenes from unconstrained internet photo collections of cultural landmarks, that are also unbounded, but additionally naturally plagued by illumination variations as well as transient occluders.
@@ 2 Related Work
Subsequent to NeRFs, several techniques emerged to extend the NeRF setup to “in-the-wild” unconstrained photo collections plagued by illumination variations and transient occluders. This added variability makes learning a scene particularly challenging, as surfaces can exhibit significant visual disparities across views. NeRF-W (Martin-Brualla et al., 2021) and Ha-NeRF (Chen et al., 2022b) address the challenge of novel view synthesis in-the-wild (NVS-W) by modeling scene lighting through appearance embeddings, and transient occluders through transient embeddings. CR-NeRF (Yang et al., 2023) improves upon previous methods by leveraging interactive information across rays to mimic the perception of humans. Fridovich-Keil et al. (2023) present K-Planes, which modify and extend Tri-Planes (Chan et al., 2022) to in-the-wild scenes thanks to learnable appearance embeddings (similarly to Martin-Brualla et al. (2021)). Our work aims to extend in-the-wild scene representations (particularly K-Planes in our case) beyond closed-world setups, by leveraging pre-trained networks.
@@
The integration of pre-trained priors for downstream tasks has emerged as a prominent trend, as they enable the effective incorporation of extrinsic knowledge into diverse applications. For neural representations, priors have been utilized for few-shot scene modeling (Yu et al., 2021; Jain et al., 2021; Deng et al., 2023) as well as generative tasks (Shue et al., 2023; Poole et al., 2023; Watson et al., 2023) for object generation and novel view synthesis. In this realm, denoising diffusion probabilistic models (Ho et al., 2020; Rombach et al., 2022) have recently gained particular attention, and have seen applications as plug-and-play priors (Graikos et al., 2022) and utilized in various domains such as super-resolution (Wang et al., 2023) and more specifically novel view synthesis. Liu et al. (2023) fine-tune Stable Diffusion (Rombach et al., 2022), a pre-trained latent diffusion model for 2D images, to learn camera controls over a 3D dataset and thus performing NVS by generalizing to other objects. These results hold paramount value, as they highlight the rich 3D prior learned by Stable Diffusion, even though it has only been trained on 2D images. This however comes with geometric inconsistency issues across views, as a pre-trained model alone has no explicit multi-view geometric constraints. Hence, we look for a way to integrate pre-trained networks into scene learning that sidesteps the aforementioned problems. As presented, pre-trained networks for implicit-field representations have seen various applications. However, the exploitation of pre-trained priors for in-the-wild applications remains an unexplored area of research.
@@
3 Method
To guide the optimization of scene representations with extrinsic signals, we learn a scene through two alternating stages, as illustrated in Figure 2. Scene fitting optimizes our K-Planes representation $\mathbf{P}_{\gamma}$ to reproduce the images in the training set, as traditionally done in neural rendering techniques. Scene refining finetunes a pre-trained network to this K-Planes representation, and then infers a new one $\mathbf{P}_{\varepsilon}$, which will subsequently be corrected by scene fitting. The main idea behind this is that we use our 3D implicit model $\mathbf{P}_{\gamma}$ for optimizing the scene to the available information in the training set and adhere to essential geometric constraints, and then project this scene representation on the set of scenes inferable by the pre-trained network, making it closer to natural images. In this section, we detail each stage and elucidate the intuition behind our method.
@@ 3.1 Scene Fitting
The goal at this stage is to fit a scene, adhering to pre-defined geometric constraints, from posed RGB images. To fit the scene, we adopt the K-Planes representation (Fridovich-Keil et al., 2023) (closely related to Tri-Planes (Chan et al., 2022) and HexPlanes (Cao & Johnson, 2023)). As such, this stage corresponds to following the optimization procedure of K-Planes representations as presented in (Fridovich-Keil et al., 2023), from which we adapt the code.
K-Planes are compact 3D model representations applicable to static scenes, “in-the-wild” scenes (scenes with varying appearances), and dynamic scenes. These models allow for fast training and rendering, while maintaining low-memory usage. K-Planes model a d-dimensional scene with $k=\binom{d}{2}$ planes, which represent the combinations of every pair of dimensions. This structure makes K-Planes compatible with a multitude of neural network architectures, and more particularly image-specialized network architectures. This enables K-Planes generation by minimally tweaking generative image architectures. For a static 3D scene, $k=3$ and the planes represent the $xy,xz,$ and $yz$ planes. These planes, each of size $N\times N\times C$, encapsulate features representing the density and view-dependent colors of the scene.
@@
where $\hat{C}(\mathbf{r})$ is the expected color, $T_{i}$ is the accumulated transmittance along the ray, and $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent samples. Proposal sampling is also employed similarly to Fridovich-Keil et al. (2023) to iteratively refine a small instance of K-Planes (using histogram loss (Barron et al., 2022)) that estimates densities along a ray. These density estimates are used to allocate more points in regions of higher densities.
Figure 3: Synthetic toy example. Qualitative results on the Lego scene from the NeRF synthetic dataset (Mildenhall et al., 2020) showcasing the optimization progression on RefinedFields, and a comparison with the ground truth and K-Planes. The training set is constrained to 50% of its initial size for both RefinedFields and K-Planes. RefinedFields refines the K-Planes representation enabling the proper reconstruction of details in the scene. At the end of optimization, the Mean Squared Error (MSE) for RefinedFields is $3.46\times 10^{-4}$, while the one for K-Planes is $4.36\times 10^{-4}$.

### v4

# algorithmic

@@ Abstract
Planar scene representations have recently witnessed increased interests for modeling scenes from images, as their lightweight planar structure enables compatibility with image-based models. Notably, K-Planes have gained particular attention as they extend planar scene representations to support in-the-wild scenes, in addition to object-level scenes. However, their visual quality has recently lagged behind that of state-of-the-art techniques. To reduce this gap, we propose RefinedFields, a method that leverages pre-trained networks to refine K-Planes scene representations via optimization guidance using an alternating training procedure. We carry out extensive experiments and verify the merit of our method on synthetic data and real tourism photo collections. RefinedFields enhances rendered scenes with richer details and improves upon its base representation on the task of novel view synthesis. Our project page can be found at https://refinedfields.github.io .
@@ 1 Introduction
[RefinedFields Rendering]Refer to caption Refer to caption Refer to caption Refer to caption
@@
To draw novel objects and views, humans often rely on a blend of cognition and intuition, where the latter is built on a large prior acquired from a long-term continuous exploration of the visual world. Nevertheless, ablating one of these two elements results in catastrophic representations. On the one hand, humans find it particularly difficult to draw a bicycle based solely on this preconceived prior (Gimini, 2016). However, once one photograph is observed, drawing novel views of a bicycle becomes straightforward. On the other hand, drawing monuments and complex objects based solely on observed images, and with no preconceived notions of geometry and physics, is also non-trivial. In computer vision, recent methods tackling object generation and novel view synthesis generally focus on either the former or the latter ablation.
@@
The first class of methods approaches novel view synthesis by learning scenes through rigorous cognition, as in dense observations of captured images. Although classic methodologies like structure-from-motion (Hartley & Zisserman, 2004) and image-based rendering (Shum et al., 2008) have previously tackled this problem, the field has recently seen substantial advancements thanks to stereo reconstruction techniques (Wang et al., 2024), as well as implicit neural representation methods (Mildenhall et al., 2020) which this paper focuses on. Particularly, planar scene representations such as Tri-Planes (Chan et al., 2022) have recently gained significant attention (Shue et al., 2023; Lan et al., 2024) for their lightweight planar structure that allows for a seamless integration with image-based models. Recent methods extend NeRFs (Martin-Brualla et al., 2021) and Tri-Planes (Fridovich-Keil et al., 2023, K-Planes) to support learning from unconstrained “in-the-wild” photo collections by enabling robustness against illumination variations and transient occluders. These representations, however, do not learn any prior across scenes as they are trained from scratch for each scene. This means that these representations are learned in a closed-world setting, where the information scope is limited to the training set at hand. Despite the wide interest in planar scene representations, their visual quality currently falls short behind recent state-of-the-art methods (Kulhanek et al., 2024; Xu et al., 2024). Our goal is to integrate pre-trained networks into the training framework of K-Planes, with an aim to improve their rendering performance.
The second class of methods tackles novel view synthesis and object generation by learning and leveraging priors over images and scenes, reminiscent of drawing from insights and intuition. These methods have also recently witnessed accelerated advancements. Recent works leverage pre-trained networks to achieve Object Generation (OG) (Poole et al., 2023; Metzer et al., 2023) and Novel View Synthesis (NVS) (Yu et al., 2021; Jain et al., 2021; Liu et al., 2023; Melas-Kyriazi et al., 2023). Particularly, Liu et al. (2023) achieve NVS from single images only by simply fine-tuning a pre-trained latent diffusion model (Rombach et al., 2022). This proves pivotal significance related to large-scale pre-trained vision models, as it shows that, although trained on 2D data, these models learn a rich geometric 3D prior about the visual world. Nevertheless, as these pre-trained models alone have no explicit multi-view geometric constraints, their use for 3D applications is usually prone to geometric issues (e.g.  geometric inconsistencies, multi-face Janus problem, content drift issues (Shi et al., 2023, Figure 1)). This class of methods has not yet been explored to enhance implicit models representing in-the-wild scenes, as leveraging priors over such representations is not evident.
RefinedFields proposal. Our work builds on the previous discourse and aims to enhance planar scene modeling by leveraging pre-trained networks. To this end, we present an alternating training algorithm that integrates a novel scene refining stage, which aims to propose a better optimization initialization by utilizing a pre-trained model. We adopt K-Planes (Fridovich-Keil et al., 2023) as a base scene representation, which extend Tri-Planes to support in-the-wild scenes, in addition to object-level scenes. Liu et al. (2024) highlight that features in planar scene representations resemble projected scene images, a finding we also corroborate in Appendix B, and exploit to improve upon K-Planes. Specifically, RefinedFields refines planar scene representations by projecting them onto the space of representations inferable by a pre-trained network, which pushes K-Planes features to more closely resemble real-world images. To do so, we build on the seamless integration of K-Planes with image-based models and present an alternating training procedure that iteratively switches between optimizing a K-Planes representation on images from a particular dataset, and fine-tuning a pre-trained network to output a new conditioning leading to a refined version of this K-Planes representation. Overall, this procedure guides the optimization of a particular scene, by leveraging not only the training dataset at hand but also the rich prior lying within the weights of the pre-trained model, which is a first for in-the-wild scene modeling.
We conduct extensive quantitative and qualitative evaluations of RefinedFields. We show that our method improves upon K-Planes with richer details in scene renderings. We prove via ablation studies that this added value indeed comes from the fine-tuned prior of the pre-trained network. Figure 1 illustrates the improvements our method showcases on the Trevi fountain scene from Phototourism (Jin et al., 2020).
2 Related Work
RefinedFields achieves geometrically consistent novel view synthesis, that can also be applied in-the-wild, by leveraging K-Planes, and a large-scale pre-trained network. Our method is the first to satisfy all of these attributes, as summarized in Table 1. In this section, we develop the various preceding works from which our method takes inspiration.
Neural representations.
Neural rendering (Tewari et al., 2020) has seen significant advancements since the introduction of Neural Radiance Fields (Mildenhall et al., 2020, NeRF). At its core, NeRF learns a scene by fitting the weights of a neural network on posed images of said scene. This subsequently enables the reconstruction of the scene thanks to volume rendering (Kajiya & Herzen, 1984). Subsequent to NeRFs, various scene representations have been introduced (Chen et al., 2022a; Müller et al., 2022; Kerbl et al., 2023). Particularly, Chan et al. (2022) introduce Tri-Planes, a planar scene representation serving as a middle ground between implicit and explicit representations, enabling a faster learning of scenes. Tri-Planes have been widely adopted in recent works (Shue et al., 2023; Lan et al., 2024; Wang et al., 2023b), as their planar structure enables a seamless integration with image-based models.
@@ 2 Related Work
Table 1: Related work overview. RefinedFields leverages a pre-trained prior (Rombach et al., 2022) on the scene representation to refine K-Planes, our underlying scene representation, utilized for novel view synthesis in the wild (NVS-W).
Non exhaustive, other works with characteristics similar to NeRF-W exist. Refer to Section 2 for more details.
@@
NFD (Shue et al., 2023) Tri-Planes OG
3DGen (Gupta et al., 2023) Tri-Planes
Latent-NeRF (Metzer et al., 2023) NeRF
DreamFusion (Poole et al., 2023) NeRF
NeRF (Mildenhall et al., 2020) NeRF NVS
RealFusion (Melas-Kyriazi et al., 2023) NeRF
DiffusioNeRF (Wynn & Turmukhambetov, 2023) NeRF
NerfDiff (Gu et al., 2023) Tri-Planes
3DiM (Watson et al., 2023)
Zero-1-to-3 (Liu et al., 2023)
NeRF-W (Martin-Brualla et al., 2021) NeRF NVS-W
WildGaussians (Kulhanek et al., 2024) 3DGS
Splatfacto-W (Xu et al., 2024) 3DGS
K-Planes (Fridovich-Keil et al., 2023) K-Planes
RefinedFields (ours) K-Planes
@@
Subsequent to NeRFs, several techniques emerged to extend the NeRF setup to “in-the-wild” unconstrained photo collections (Jin et al., 2020, Phototourism) plagued by illumination variations and transient occluders. This added variability makes learning a scene particularly challenging, as surfaces can exhibit significant visual disparities across views. NeRF-W (Martin-Brualla et al., 2021) addresses the challenge of novel view synthesis in-the-wild (NVS-W) by modeling scene lighting through appearance embeddings, and transient occluders through an additional transient head. Splatfacto-W (Xu et al., 2024) proposes a Nerfstudio (Tancik et al., 2023) implementation of 3D Gaussian Splatting (Kerbl et al., 2023, 3DGS) that extends standard 3DGS to support in-the-wild scene modeling via appearance embeddings. WildGaussians (Kulhanek et al., 2024) also proposes using 3DGS with appearance embeddings for in-the-wild scene modeling, but additionally predicts uncertainty masks to exclude transient occluders from the loss computation. It is important to note that although this work adopts a pre-trained network, it is solely used for the computation of the loss masks on the data, and not as a prior on the scene representation. This approach is orthogonal to our approach of leveraging a prior to directly improve upon a specific scene representations, as it acts on the data and not the scene representation itself. Fridovich-Keil et al. (2023) present K-Planes, which modify and extend Tri-Planes (Chan et al., 2022) to in-the-wild scenes thanks to learnable appearance embeddings, similarly to Martin-Brualla et al. (2021). Our work aims to improve upon K-Planes representations by extending their training beyond closed-world setups using pre-trained networks. Note that other works also tackle NVS-W: Chen et al. (2022b) model scene lighting through appearance embeddings and transient occluders through transient embeddings, Yang et al. (2023) leverage interactive information across rays to mimic the perception of humans in-the-wild, and Chen et al. (2024) separate static and transient components by utilizing heuristics-guided segmentation. However, these works limit their training and evaluations to downscaled versions of Phototourism, which makes their results not directly comparable with ours.
@@
The integration of pre-trained priors for downstream tasks has emerged as a prominent trend, as they enable the effective incorporation of extrinsic knowledge into diverse applications. For neural representations, priors have been utilized for few-shot scene modeling (Yu et al., 2021; Jain et al., 2021; Deng et al., 2023; Gao et al., 2024) as well as generative tasks (Shue et al., 2023; Poole et al., 2023; Watson et al., 2023; Gao et al., 2024) for object generation and novel view synthesis. In this realm, denoising diffusion probabilistic models (Ho et al., 2020; Rombach et al., 2022) have recently gained particular attention for their application as plug-and-play priors (Graikos et al., 2022). They have been utilized in various domains such as super-resolution (Wang et al., 2023a) and more specifically novel view synthesis. Liu et al. (2023) fine-tune Stable Diffusion (Rombach et al., 2022), a pre-trained latent diffusion model for 2D images, to learn camera controls over a 3D dataset and thus performing NVS by generalizing to other objects. These results hold paramount value, as they highlight the rich 3D prior learned by Stable Diffusion, even though it has only been trained on 2D images. This however comes with geometric inconsistency issues across views, as a pre-trained model alone has no explicit multi-view geometric constraints. In this work, we aim to leverage a pre-trained model to enhance a volumetric scene representation, which inherently adheres to consistency constraints, hence mitigating multi-view consistency issues. The utilization of such priors to enhance in-the-wild scene representation has been an unexplored area of research.
3 Method
To guide the optimization of planar scene representations with extrinsic signals, we learn a scene through two alternating stages, as illustrated in Figure 2. Scene fitting optimizes our K-Planes representation $\mathbf{P}_{\gamma}$ to reproduce the images in the training set, as traditionally done in neural rendering techniques. Scene refining finetunes a pre-trained network to this K-Planes representation, and then infers a new one $\mathbf{P}_{\varepsilon}$, which will subsequently be corrected by scene fitting. The main idea behind this is that we use our 3D implicit model $\mathbf{P}_{\gamma}$ for optimizing the scene to the available information in the training set and adhere to essential geometric constraints, and then project this scene representation on the set of scenes inferable by the pre-trained network, making it closer to natural images. In this section, we detail each stage and elucidate the intuition behind our method.
@@ 3.1 Scene Fitting
The goal at this stage is to fit a scene, adhering to pre-defined geometric constraints, from posed RGB images. To fit the scene, we adopt the K-Planes representation (Fridovich-Keil et al., 2023). As such, this stage corresponds to optimizing K-Planes to fit a scene, from which we adapt the code.
K-Planes are compact 3D model representations applicable to static scenes, “in-the-wild” scenes (scenes with varying appearances), and dynamic scenes. These models allow for fast training and rendering, while maintaining low-memory usage. K-Planes model a d-dimensional scene with $k=\binom{d}{2}$ planes, which represent the combinations of every pair of dimensions. This structure makes K-Planes compatible with a multitude of neural network architectures, and more particularly image-specialized network architectures. This enables K-Planes inference by minimally tweaking image architectures. For a static 3D scene, $k=3$ and the planes represent the $xy,xz,$ and $yz$ planes. These planes, each of size $N\times N\times C$, encapsulate features representing the density and view-dependent colors of the scene.
@@
where $\hat{C}(\mathbf{r})$ is the expected color, $T_{i}$ is the accumulated transmittance along the ray, and $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent samples.